The near point of a patient's eye is 50.0cm. What focal length must a corrective lens have to enable the eye to see clearly an object 25cm away? Neglect the eye-lens distance

1
Expert's answer
2022-03-15T10:34:08-0400

Given:

"d_i=-50.0\:\rm cm"

"d_o=25.0\:\rm cm"

The thin lens equation says

"\frac{1}{f}=\frac{1}{d_0}+\frac{1}{d_i}"

So

"\frac{1}{f}=\frac{1}{25.0}+\frac{1}{-50.0}=\frac{1}{50.0}"

"f=50\:\rm cm"